rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_green
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#8 — FS#3944 Switch replacements in rooms 1, 2 and 3

Attached to Project— Dedicated servers
Maintenance
all (dedicated servers)
CLOSED
100%
We start a series of switches replacement in our data center in Paris 19.
The entire park will be replaced by newer models with new features.

The replacement will be performed at night, causing a break of 2-5 minutes on each server.



Date:  Monday, 06 September 2010, 12:58PM
Reason for closing:  Done
Comment by OVH - Tuesday, 11 May 2010, 16:03PM

This morning, we are carrying out the replacement of the switch of the A01 bay.
The next ones will be done at night.